Given utility function U=X0.5Y0.5 Where Px=4 Birr and py=4 Birr and the income of the consumer M=240

a. Find the utility maximizing combination of X and Y.

b. Calculate marginal rate of substitution of X for Y (MRSx,y) at equilibrium and interpret your result


1
Expert's answer
2021-12-27T08:59:33-0500

Solution:


A)



"MU_x=0.5 \frac {y^{0.5}}{x^{0.5}}""MU_y=0.5 \frac {x^{0.5}}{y^{0.5}}""{\frac{MU_x}{p_x}}=\frac {MU_y}{p_y}""x \times p_x+ y \times p_y=M""y=3x""4 \times x+4\times y=240""x=15, y=45."



D)



"\frac {\partial U}{\partial x \partial y} =\frac {0.25} {(xy)^{0.5}}""MRS x.y = \frac {\partial U} {\partial x \partial y} = \frac {0.25}{(15 \times 45)^{0.5}}=0.00962"